OpenStudy (anonymous):

Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 Could someone help me figure this out please?

OpenStudy (anonymous):

User: Marshall Davis In Course: Algebra I V20 ( 4077) Instructor: Donna Schroeder Exam: 05.04 Module Five Quiz Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Question 1 (Multiple Choice Worth 5 points) (05.01 LC) Systems of equations with different slopes and different y-intercepts have more than one solution. Always Sometimes Never Question 2 (Multiple Choice Worth 5 points) (05.02 MC) What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 −1 0 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system Question 3 (Multiple Choice Worth 5 points) (05.02 MC) Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 He should have substituted 5 + 2x He did not distribute correctly He combined like terms incorrectly, it should have been 4x instead of 5x He made no mistake That is with the answer choices, my bad.